OpenStudy (anonymous):

simplify -2 square root of 10 × square root of 8

OpenStudy (coconutjj):

\[-2\sqrt{10}\times \sqrt{8}\]

OpenStudy (coconutjj):

\[-8\sqrt{5}\]